A plate 100 mm wide, 200 mm long, and 12 mm thick is loaded in tension in the direction of the length. The plate contains an edge crack with the crack length of 16 mm. The material is steel with KIC = 80 MPa√m, and Sy = 950 MPa. Determine the maximum possibleload that can be applied before the plate (a) yields, and (b) has uncontrollable crack growth.


1
Expert's answer
2022-02-06T11:35:09-0500

load that will cause uncontrollable crack growth.

"P =\frac{K_{IC}\times A}{\beta\times \sqrt{\pi a}} =\frac{80\times 1200 }{1.3\times \sqrt{\pi 12}} =12.03" kN `